solving 2x^2=7x-29
1
Expert's answer
2019-12-04T09:44:17-0500

"2x^2 -" 7x +29=0,

a= 2, b=-7, c=29,

D= b2-4ac,

D=(-7)2-4*2*29=49-232=-183<0, hence the equation has no real solutions.